Dearest tutors, can you help me regarding my problem? thank you for that..
1. A chemist wanted to mix milk testing 3% butterfat with cream and testing 28% butterfat. How many quartz of each kind must the chemist use to produce a mixture of 10 gallons which shall test 18% butterfat?

Answer by ptaylor(2198)

Let x= Number of qts of milk (3% butterfat)
Then 40-x=Number of qts of cream (28% butterfat)
(Note: We need to deal either in qts or gallons. I have chosen to deal in qts and 10 gal=40 qts)
Now we know that the amount of pure butterfat in the milk (x)(.03) plus the amount of pure butterfat in the cream (40-x)(.28) must equal the amount of pure butterfat in the final mixture (40)(.18). Thus, our equation to solve is:
(x)(.03)+(40-x)(.28)=(40)(.18) simplifying, we get:
.03x+11.2-.28x=7.2 or
-.25x=-4
x=16 qts of milk
40-x=40-16=24 qts of cream
Hope this helps ------ptaylor